# Neuro Symbolic Knowledge Reasoning for Procedural Video Question Answering with Knowledge Module Learning

### v2

@@ Abstract
We introduce PKR-QA (Procedural Knowledge Reasoning Question Answering), a new benchmark for question answering over procedural tasks that require structured reasoning. PKR-QA is constructed semi-automatically using a procedural knowledge graph (PKG), which encodes task-specific knowledge across diverse domains. The PKG is built by curating and linking information from the COIN instructional video dataset and the ontology, enriched with commonsense knowledge from ConceptNet and structured outputs from Large Language Models (LLMs), followed by manual verification. To generate question-answer pairs, we design graph traversal templates where each template is applied systematically over PKG. To enable interpretable reasoning, we propose a neurosymbolic approach called Knowledge Module Learning (KML), which learns procedural relations via neural modules and composes them for structured reasoning with LLMs. Experiments demonstrate that this paradigm improves reasoning performance on PKR-QA and enables step-by-step reasoning traces that facilitate interpretability. Our theoretical analysis on KML learning shows that our trained models satisfy near optimal conditions for learning KG relations as neural network mapping models. Code and dataset will be released soon.
@@ 5.2 Analysis and Discussion
5.3 Theoretical analysis on the KML learning.
The objective of KML is to learn a neural network mapping function
$\phi_{{R}_{k}}(~x(e_{i};~\theta_{x});~\theta_{{R}_{k}})\rightarrow\mathbf{e_{j}}$
where $\mathbf{e_{j}}$ is a vector that is closer to $x(e_{j})$ for all $e_{j}\in\mathcal{I}_{j}$.
For this theoretical analysis without loss of generality let us introduce the following problem to study neural mapping such as $\phi_{{R}_{k}}$.
Let $X=\{x_{1},x_{2},\dots,x_{n}\}$ be a finite set of vectors representing the learned embeddings of each entity in the KG. For each $x\in X$ we define:
  • $Y(x)\subseteq X$: the set of positive target entity embeddings for $x$ under the relation.
  • $U(x):=X\setminus Y(x)$: the set of negatives entity embeddings under the relation.
  • $z=\phi(x)$: the vector mapping of $x$ under the relation.
  • Each $z:=\frac{z}{\|z\|_{2}}$: the normalized embedding (unit vector).
Then let us define cosine similarity
$s(a,b):=\langle a,b\rangle,\qquad a,b\in\mathbb{R}^{D},\;\|a\|=\|b\|=1,$
so that $s(a,b)\in[-1,1]$.
Similarity-based constraint for KML optimization
The optimal condition for KML learning is to make sure that
$\min_{y\in Y(x)}s\!\left(\phi(x),y\right)\;>\;\max_{u\in U(x)}s\!\left(\phi(x),u\right).$ (9)
. The cosine-similarity-based constraint requires that the least similar positive still has higher cosine similarity (with $z$) than the most similar negative.
Implications for the mapping $\phi$
The above constraints imply several structural requirements on $\phi$:
  • Unit-normalization: Outputs must be normalized:
    $\|\phi(x)\|_{2}=1,\quad\forall x\in X.$
  • Non-collapse: $\phi$ must not map all inputs to the same unit vector, otherwise eq. 9 cannot be satisfied.
  • Continuity: $\phi$ should be Lipschitz continuous in the angular metric, so that small perturbations in $x$ do not drastically change $\phi(x)$.
Bound on the contrastive loss
We define the multi-positive contrastive loss for KML as follows:
$\mathcal{L}(x)\;=\;-\log\frac{\sum\limits_{y\in Y(x)}\exp\!\big{(}s(\phi(x),y)/\tau\big{)}}{\sum\limits_{v\in X}\exp\!\big{(}s(\phi(x),v)/\tau\big{)}},$ (10)
where $\tau>0$ is a temperature parameter. For simplicity of the proof we drop the temperature.
Let us denote $s_{v}:=s(\phi(x),v)$. Then let us define A, B and C as follows:
$A=\sum_{y\in Y(x)}e^{s_{y}},\quad C=\sum_{u\in U(x)}e^{s_{u}},\quad B=A+C,\quad$
Let us denote $\varepsilon$ to be such that $\mathcal{L}(x)\leq\varepsilon$. Then we obtain
$\frac{A}{A+C}\;\geq\;e^{-\varepsilon}\;\;\Longleftrightarrow\;\;C\;\leq\;(e^{\varepsilon}-1)\,A.$
We know that
$\sum_{i=1}^{|Y(x)|}e^{\min_{y\in Y(x)}s_{y}}\leq\sum_{y\in Y(x)}e^{s_{y}}$
$\Rightarrow|Y(x)|e^{\min_{y\in Y(x)}s_{y}}\leq A\\$
$\Rightarrow\min_{y\in Y(x)}s_{y}\leq log(A)-log(|Y(x)|)$
Similarly,
$\sum_{u\in U(x)}e^{s_{u}}\geq e^{\max_{u\in U(x)}s_{u}}$
$\Rightarrow\max_{u\in U(x)}s_{u}\leq log(C)$
From (i) and (iii)
$\max_{u\in U(x)}s_{u}\leq log((e^{\varepsilon}-1)\,A)$
$\Rightarrow\max_{u\in U(x)}s_{u}\leq log(e^{\varepsilon}-1)+log(A)$
From (ii) - (iv)
$\min_{y\in Y(x)}s_{y}-\max_{u\in U(x)}s_{u}\leq-log(|Y(x)|)-log(e^{\varepsilon}-1)$
$\Rightarrow\min_{y\in Y(x)}s_{y}-\max_{u\in U(x)}s_{u}\leq\log\!\Bigg{(}\frac{1}{|Y(x)|\,(e^{\varepsilon}-1)}\Bigg{)}.$
As the left hand size should be >0>0, the $\log\!\Bigg{(}\frac{1}{|Y(x)|\,(e^{\varepsilon}-1)}\Bigg{)}$ can be >0>0 iff $\frac{1}{|Y(x)|\,(e^{\varepsilon}-1)}>1$. Therefore,
$1+Y(x)>Y(x)e^{\varepsilon}$
$\Rightarrow log(1+Y(x))-log(Y(x))>\varepsilon$
$\Rightarrow\varepsilon<log(1+Y(x))-log(Y(x))$
$\Rightarrow\varepsilon<log(1+\frac{1}{Y(x)})$
As $L(x)\leq\varepsilon$, if
$\mathcal{L}(x)<\log\!\left(1+\frac{1}{|Y(x)|}\right),$ (10)
then the right-hand side is positive, which implies
$\min_{y\in Y(x)}s_{y}\;>\;\max_{u\in U(x)}s_{u},$
and therefore cosine constraint holds.
In our experiments, we obtain a loss of 0.71 which aligns with this bound. The temperature simply rescales the needed loss threshold by $\tau$(larger $\tau$ makes the separation easier but gradients smaller). Therefore, $\tau$ has to be set depending on the batch size. We set the $\tau$ to be batch size.
As our KML neural network can be written as follows:
$f(x)=h(g(x))$
where
$h(x)=\frac{x}{\|x\|}$
and
$g(x)=W_{1}tanh(W_{2}x).$
As
[…]
and
$g(u)-g(v)||\leq\|W_{1}\|\cdot\|W_{2}\|$
The Lipschitz constant for our Knowledge Modules is given by
$L=\frac{2}{\delta}\,\|W_{1}\|\cdot\|W_{2}\|,$
where $\delta$ denotes the minimum norm of $g(x)$ over all $x$. This implies that, as long as the neural network $g(x)$ does not produce vectors with very small norms, learning can proceed effectively.
In particular, this result shows that our neural networks are Lipschitz continuous whenever $\delta>0$. It also motivates the use of weight regularization (e.g., weight decay), which we apply with a decay factor of $0.01$ during training.
Moreover, this analysis suggests that when learning embeddings, it may be beneficial to use larger embedding dimensions $D$, as they help avoid excessively small vector norms and thus improve stability. The choice of $W_{1}\in\mathcal{R}^{d\times D}$ and $W_{2}\in\mathcal{R}^{d\times D}$ may also improve stability.

### v3

@@ Abstract
We introduce PKR-QA (Procedural Knowledge Reasoning Question Answering), a new benchmark for question answering over procedural tasks that require structured reasoning. PKR-QA is constructed semi-automatically using a procedural knowledge graph (PKG), which encodes task-specific knowledge across diverse domains. The PKG is built by curating and linking information from the COIN instructional video dataset and the ontology, enriched with commonsense knowledge from ConceptNet and structured outputs from Large Language Models (LLMs), followed by manual verification. To generate question-answer pairs, we design graph traversal templates where each template is applied systematically over PKG. To enable interpretable reasoning, we propose a neurosymbolic approach called Knowledge Module Learning (KML), which learns procedural relations via neural modules and composes them for structured reasoning with LLMs. Experiments demonstrate that this paradigm improves reasoning performance on PKR-QA and enables step-by-step reasoning traces that facilitate interpretability. Code and dataset will be released soon https://github.com/LUNAProject22/KML.